Parikshit makes a cuboid of plasticine of sides 5 cm, 2 cm, 5 cm. How many such cuboids will he need to form a cube?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


side of the cube is 5 cm, 2 cm and 5 cm.

Therefore, volume of cube = 5×2×5 = 50

The prime factorisation of 50 = 2×5×5

Here, 2, 5 and 5 cannot be grouped into triples of equal factors.

Therefore, we will multiply 50 by 2×2×5 = 20 to get perfect square.

Hence, 20 cuboid is needed.

Thanks